OpenStudy (anonymous):

A shirt cost $10 and a skirt cost twice as much. Nicole bought 2 more shirts than skirts. She spent $80 altogether. Hoe many blouses did she buy?

OpenStudy (anonymous):

Well, what we know is that blouses (shirts) are $10. Therefore, if skirts cost twice as much they would be $20. Right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Okay. So it also says that she bought at least two shirts. So that right there is $20 and she also bought at least 1 skirt so that is $40.

OpenStudy (anonymous):

One sec. Trying to put it in a better way.. One that you will understand and learn from. Lol

OpenStudy (anonymous):

It's fine, thank you

OpenStudy (anonymous):

She bought 4 blouses and two skirts. 10x4=40+2x20=40. 40+40=80

OpenStudy (anonymous):

Thank you! Good explanation